The population of a town grows at a rate proportional to the population present at time t. The initial population of 500 increases by 15% in 10 years. What will be the population in 50 years? (Round your answer to the nearest person.)

____ persons

How fast is the population growing at t = 50? (Round your answer to two decimal places.)

______ persons / year
1
Expert's answer
2020-06-22T15:35:01-0400

According to the question,

"\frac{dP}{dt} = kP" where P is population at any time


solving equation,

"\int \frac{dP}{P} = \int kdt"


"lnP = kt + C" (1)

Applying the conditions,

at t=0, P = 500


"ln500 = C" (2)


at t=10, P= 575


"ln(575) = k(10) + ln(500)"

"k = \frac{1}{10} ln (\frac{575}{500}) = 0.0139762"


So Equation of the Population growth is given by


"P = 500 e^{0.0139762t}" (3)



Population in 50 years,

"P = 500 e^{0.0139762*50} = 1005.6 \approx 1006" people



Rate of population growth is

"\frac {dP}{dt} = \frac {d (500 e^{0.0139762t})}{dt} = (500*0.0139762)e^{0.0139762t}"

putting t = 50, we get


"\frac {dP}{dt} =(500*0.0139762) e^{0.0139762*50} = 14" people/year (approx)